# Frostman random variables, entropy inequalities, and applications

### v4

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum-product phenomena in both independent and dependent settings. Fix $0<s<1$, and let $(X,Y)$ be a bivariate real random variable with bounded support, whose distribution satisfies a Frostman condition of dimension $s$. Let $\phi(x,y)$ be a polynomial obtained from a diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)\in\mathbb{R}[x,y]$ of degree $d\geq 2$ by applying an invertible rational linear change of variables in $(x,y)$. We show that there exists $\epsilon=\epsilon(\phi,s)>0$ such that
@@
for all sufficiently large $n$, where the precise assumptions on $(X,Y)$ depend on the Frostman level. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
As an application, we obtain discretized sum-product type estimates. In particular, for a $\delta$-separated set $A\subseteq[0,1]$ of cardinality $\delta^{-s}$, satisfying certain non-concentration conditions, and a dense subset $G\subseteq A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
@@ 1 Introduction
More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e. $a_{2}^{2}-4a_{1}a_{3}\neq 0$, and $(X,Y)$ is a bivariate Frostman random variable with bounded range.
For $\phi(x,y)=x\cdot y$, we recall the following theorem due to Máthé and O’Regan in [12]. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $c_{1},c_{1}]$, with $0<s<1$, and $C,c_{1}\geq 1$. Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c_{1})$ such that for all $n>N,$ we have
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, these conditions form a hierarchy:
@@
Remarkably, different levels of this hierarchy yield entropy inequalities with different strengths and applications.
@@
Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C,c>0$. There exist a positive constant $\epsilon=\epsilon(d,s)$ and an integer $N=N(\rho_{1},\rho_{2},s,C,c)$ such that the following statement holds. For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon)$
@@
In the theorem above, a constant depends on a polynomial $\rho$, we means that it depends on the degree and coefficients of $\rho$. The notation $1\ll_{s}s_{1},s_{2}$ means that $s_{1}$ and $s_{2}$ are bounded below by some positive constant depending only on $s$, but not on $\rho_{1},\rho_{2},C,c$ or $n$.
@@
Note that $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. We say that $\phi(x,y)$ is diagonal if exactly two of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ are zero. In such a situation, one has $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ for some polynomials $\tilde{\rho}_{1}(x)$ and $\tilde{\rho}_{2}(y)$ of degree $d$. And we say that $\phi(x,y)$ is non-diagonal if at most one of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ is zero.
@@
Let $\phi$ be given as in (1). Let $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(d,s)$, $\epsilon^{\prime}=\epsilon^{\prime}(d,\Xi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$
Here, a constant depends on the matrix $\Xi$ means that it depends on the entries of $\Xi$. And a constant depends on $\phi$ means it depends on $\rho_{1}$, $\rho_{2}$ and $\Xi$.
Theorem 1.4 (Marginally Frostman with independence).
Let $\phi$ be given as in (1). Let $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $X$ and $Y$ be two independent real random variables whose ranges are contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$ (2)
@@
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 7.1 for an example matching this bound). We show in Examples 7.1 and 7.2 that there are i.i.d. random variables $X$ and $Y$ such that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
@@
    The positive $\epsilon$ might fail to exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
  3. 3.
@@
    as $n\rightarrow\infty$, where $Y$ is i.i.d. with $X$. In other words, the term $ns$ cannot be replaced by $H_{n}(X)$ in the right-hand side of (2).
@@
    In Theorems 1.2 and 1.3, the necessity of different terms in these maxima (for example, $\frac{H_{n}(X,Y)}{2}$ or $\frac{H_{n}(X)+H_{n}(Y)}{2}$) reflects subtle aspects of how dependency affects entropy, we provide an example showing it cannot be simplified in the statement of Theorem 1.2 (Example 7.5).
@@
    In the statement of Theorem 1.4, the fact that NN depends on $C$, and the condition that $X$ and $Y$ are independent are both essential (see Examples 7.4 and 7.5, respectively).
  6. 6.
    The value of $\epsilon$ is not explicit, and our arguments suggest that they are quantitatively small.
We record several further remarks. Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e. $H_{n}\big((X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big)$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e. $\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function. By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.4 for the full range of $s$ for this specific bilinear form function. We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$. Moreover, in terms of applications, i.e. sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors. It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
We now discuss applications of the entropy theorems.
For $A\subseteq\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
@@
By Theorem 1.3, extending this result to more general polynomials is possible, but requiring minimum degree constraints (3) and (4).
@@
In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subseteq A$ with $B\geq\delta^{\varepsilon}(\#A)$.

### v5

# arXiv:2507.15196v5[math.CA] 02 Mar 2026

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum-product phenomena in both independent and dependent settings. Fix $0<s<1$, and let $(X,Y)$ be a bivariate real random variable with bounded support, whose distribution satisfies a Frostman condition of dimension $s$. Let $\phi(x,y)$ be a polynomial obtained from a diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)\in\mathbb{R}[x,y]$ of degree $d\geq 2$ by applying a change of variables $\Xi\in GL_{2}(\mathbb{Q})$ in $(x,y)$. We show that there exists $\epsilon=\epsilon(d,\Xi,s)>0$ such that
@@
for all sufficiently large $n$, where the precise assumptions on $(X,Y)$ depend on the Frostman level. The proof introduces a novel multi‑step entropy framework, combining the state-of-the-art results on the Falconer distance problem, a discretized entropy Balog–Szemerédi–Gowers mechanism, and new entropy inequalities adapted to dependent variables, to reduce general polynomials of arbitrary degree to a diagonal quadratic case. As applications, we obtain innovative discretized sum-product type estimates along dense graphs. In particular, for a $\delta$-separated set $A\subseteq[0,1]$ of cardinality $\delta^{-s}$, satisfying certain non-concentration conditions, and a dense subset $G\subseteq A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
@@ 1 Introduction
More precisely, we seek a quantitative lower bound for
$\max\bigl\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\bigr\},$
where $(X,Y)$ is a bivariate random variable satisfying one of the Frostman-type nonconcentration hypotheses introduced in Definition 1.1, and $\phi:\mathbb{R}^{2}\to\mathbb{R}$ belongs to a broad class of polynomial forms of degree $d\geq 2$ (described below) obtained from a diagonal expression $\rho_{1}(x)+\rho_{2}(y)$ by composing with a rational linear change of variables. In particular, this class contains the non-degenerate rational quadratic forms.
For $\phi(x,y)=xy$, we recall the following theorem due to Máthé and O’Regan in [19]. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in [c,c][-c,c], with $0<s<1$, and $C,c\geq 1$. Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c)$ such that for all $n>N,$ we have
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
While the usual $(s,C)$-Frostman condition for measures on $\mathbb{R}$ is standard, the three bivariate variants introduced in Definition 1.1 (joint/marginal/conditional) and the way they are used as a hierarchy in an entropy-expansion framework are, to the best of our knowledge, not part of the standard entropy sum-product toolkit. They are introduced here to separate non-concentration from dependence: different levels of the hierarchy genuinely lead to different entropy inequalities and different applications. At a common exponent level $(s_{1},s_{2})$, these conditions form the following chain:
@@
1.1 Main results on entropy estimates
@@
Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C,c>0$. There exists a positive constant $\epsilon=\epsilon(d,s)$ and an integer $N=N(\rho_{1},\rho_{2},s,C,c)$ such that the following statement holds. For any bivariate real random variable $(X,Y)$ that is jointly $(s,s,C)$-Frostman and has range contained in $c,c]^{2}$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon)$ (1)
@@
Note that $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. We say that $\phi(x,y)$ is diagonal if exactly two of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ are zero. In such a situation, one has $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ for some polynomials $\tilde{\rho}_{1}(x)$ and $\tilde{\rho}_{2}(y)$ of degree $d$. We say that $\phi(x,y)$ is non-diagonal if at most one of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ is zero.
@@
Let $\phi$ be given as in (2), $0<s<1$ and $C,c>0$. There exist positive constants $\epsilon=\epsilon(d,s)$, $\epsilon^{\prime}=\epsilon^{\prime}(d,\Xi,s)$, and an integer $N=N(\phi,s,C,c)$ such that the following statement holds. Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s,C;s,C)$-Frostman. Then
$\max\left\{H_{n}(X+Y),\,H_{n}\big(\phi(X,Y)\big)\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}\end{cases}$ (3)
for all $n>N$.
@@
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 8.1 for an example matching this bound). We show in Examples 8.1 and 8.2 that there are i.i.d. random variables $X$ and $Y$ such that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
@@
    as $n\rightarrow\infty$, where $Y$ is i.i.d. with $X$. In other words, the term nsns cannot be replaced by $H_{n}(X)$ in the right-hand side of (1) and (3).
  3. 3.
    Examples 8.1, 8.3, 8.4 and 8.5 in Section 8 show that the hypotheses in Theorem 1.3 are essentially sharp. In particular, if one of the conditions is dropped, then a positive uniform gain may no longer hold.
@@
    In Theorems 1.2, the necessity of different term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ in the maxima reflects subtle aspects of how dependency affects entropy. Example 8.5 shows that this term cannot be discarded.
@@
    We do not attempt to optimize $\epsilon$ or $\epsilon^{\prime}$. The argument produces a positive gain depending on the parameters of the problem.
1.2 New discretized sum-product type estimates along dense graphs
We now discuss some novel applications of the entropy theorems.
For a finite set $A\subseteq\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$. We say $A$ is $\delta$-separated if the distance between any two distinct elements in $A$ is at least $\delta$.
@@ 1.2 New discretized sum-product type estimates along dense graphs
Theorem 1.5 extends this statement to more general polynomials under the natural minimum-degree conditions (4) and (5).
@@
In contrast to Theorem 1.5, the next result holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subseteq A$ with $B\geq\delta^{\varepsilon}(\#A)$.